It takes 32 minutes for 3 people to paint 4 walls. How long does it talk 4 to paint 7 walls
Answer by ankor@dixie-net.com(22740)

It takes 32 minutes for 3 people to paint 4 walls.
32 * 3 = 96 people-min for 4 walls
then
96/4 = 24 people-min per wall
How long does it take 4 to paint 7 walls.
let m = no. of minutes for 4 people to paint 7 walls
4m = 7(24)
4m = 168
m = 168/4
m = 42 minutes